An unknown salt, M2Z, has a Ksp of 3.3 x 10-9. Calculate the solubility in mol/L of M2Z.

a. 2.9 x 10-5 M

b. 5.7 x 10-5 M

c. 9.4 x 10-5 M

d. 3.7 x 10-5 M
Answer #1

MTI can ☺ 10 2+ ! When a los [mt]} [2]. kp = 2 = 453 =(2512. . 453 . 453 = 3.3410-9 as? = 0.825810-9 iisa 9:4 X1014 L solubil